How many formula units make up 13.0 g of magnesium chloride (MgCl2)?
1
Expert's answer
2019-03-11T08:14:29-0400

"N_{f.u.}=\frac{m}{M} \cdot N_{A}"

"N_{f.u.}" - number of formula units

"m" - mass

"M" - molar mass

"N_{A} \approx 6.02 \cdot 10^{23} \; mol^{-1}" - Avogadro's number

"M(MgCl_{2})=1 \cdot 24.305+2 \cdot 35.453=95.211 \; g/mol"

"N_{f.u.}=\frac{13}{95.211} \cdot 6.02 \cdot 10^{23} \approx 8.22 \cdot 10^{22}"